Circular No. 8750

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2006 S2 (LINEAR) An apparently asteroidal object discovered by the LINEAR project,

and posted on the ‘NEO Confirmation Page’, has been found by J. Young to show a round coma with an outer faint shell 10′′ in diameter and a central condensation 5′′ in diameter, with no hint of tail, on CCD images taken on Sept. 18.12–18.16 UT with the Table Mountain 0.61-m f/16 reflector; his images from Sept. 19.14–19.17 show a round, diffuse coma of diameter 8′′, and again no tail seen in long exposures.

2006 UT  α_{2000}  δ_{2000}  Mag.
Sept.17.17360  18^{h}16^{m}38.^{s}52  +46^{o}40′51′′.1  19.0

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2006-S26.

T = 2007 May 11.772 TT  ω = 168.^{o}186
Ω = 114.262 }2000.0
q = 3.10570 AU  i = 98.828

SUPERNOVAE 2006fo–2006gh Two apparent supernovae have been reported following discovery on

unfiltered CCD images: 2006fp by T. Puckett and R. Gagliano (cf. IAUC 8741); and 2006gc by R. Mostardi and W. Li (LOSS/KAIT; cf. IAUC 8744). Also, an additional 18 supernovae (2006fo, 2006fq–2006gb, and 2006gd– 2006gh — most being fainter than magnitude g = 20) have been reported by J. Frieman on behalf of the Sloan Digital Sky Survey II collaboration, with details published on CBETs 624, 627, and 629; the brightest, 2006fo (a type-Ia supernova ≈ 1 week past maximum), is included below.

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006fo Sept. 16  2^{h}32^{m}38.^{s}89 + 0^{o}37′03′′.0 18.2 6′′.0 W, 0′′.6 N
2006fp Sept. 17.15 22 45 41.13 +73 09 47.8 17.7 13′′.9 E, 3′′.6 N
2006gc Sept. 18.52 4 58 50.81 + 6 58 28.6 18.2 13′′.1 E, 10′′.9 S

Additional magnitudes for 2006fp in UGC 12182: 2005 July 14, [19.6 (Puckett); 2006 Sept. 18.27 UT, 17.6 (J. Grcevich, 2.4-m Hiltner telescope, MDM Observatory). Additional KAIT magnitudes for 2006gc in UGC 3219: Feb. 11.23, [19.5; Sept. 8.53, 18.1.

2006 September 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green